OpenStudy (anonymous):

Is (x + 4) one of the factors of 3x2 + 4x – 4 ?

OpenStudy (precal):

do know how to factor or do synthetic division?

Parth (parthkohli):

Easy way to do this is to use the factor theorem. That's very easy.

Parth (parthkohli):

If x - k is a factor, then k is a zero of the function.

OpenStudy (anonymous):

no it is not 6+4x-4=2+4x and x+4 cant be its factor

OpenStudy (precal):

yes, unless teacher had other intentions

Parth (parthkohli):

So, -4 must be a zero of the function.

OpenStudy (anonymous):

Just put x = -4 in the given expression and check whether its value is becoming 0 or not..

Parth (parthkohli):

\( \color{Black}{\Rightarrow 3(-4)^2 + 4(-4) - 4 = 0?}\)

OpenStudy (anonymous):

Oh, wait, are the factors: (3x - 2) and (x - 2)?

OpenStudy (anonymous):

yes that is right

OpenStudy (precal):

 known as the slide and divide method of factoring